# Conformal Prediction with Corrupted Labels: Uncertain Imputation and Robust Re-weighting

### v2

@@ 3.1.2 Is PCP robust to inaccurate weights?
The following robustness analysis of PCP is divided into two parts. First, we consider a case where the inaccurate weights $\{\tilde{w}_{i}\}_{i=1}^{n}$ are shifted by a constant error $\delta\in\mathbb{R}$ from the true weights for all samples. In the second, we extend the analysis to a more general setting where the error varies across samples. We note that the theory developed in this section also applies to WCP, as detailed in Appendix A.1. We begin by examining the case of constant error, formulated as:
@@
We remark that, in this analysis, we do not consider the sign of the weights and allow them to become negative. We denote the sum of the true weights by ${W}_{k}:=\sum_{j=1}^{k}w_{k}$ and recall that $Q^{\texttt{WCP}}$ from (5) is the threshold constructed by WCP using the true weights. We also denote by $Q^{\texttt{CP}}$ the threshold generated by Naive CP, which is CP from Section 2.1 applied using only the uncorrupted data; see Appendix A.1.1 for more details. These notations set the ground for the conditions required for PCP to achieve the desired coverage rate.
@@
Suppose that the assumptions of Theorem 1 hold. Further, suppose that at least one of the following holds: (1) $\mathbb{P}\left(Q^{\texttt{CP}}>Q^{\texttt{WCP}}\right)\geq 1-\varepsilon$ and $\delta\geq 0$; (2) $\mathbb{P}\left(Q^{\texttt{CP}}>Q^{\texttt{WCP}},\delta<-\frac{{W}_{n+1}}{n+1}\right)\geq 1-\varepsilon$; (3) $\mathbb{P}\left(Q^{\texttt{CP}}<Q^{\texttt{WCP}},\delta>-\frac{{W}_{n+1}}{n+1}\right)\geq 1-\varepsilon$ and $\delta\leq 0$; (4) $\mathbb{P}\left(Q^{\texttt{CP}}=Q^{\texttt{WCP}}\right)\geq 1-\varepsilon$. Then, the prediction set ${C}^{\texttt{PCP}}(X^{\text{test}})$ constructed by PCP with weights shifted by $\delta$, as in (7), satisfies:
@@
We demonstrate Theorem 2 on two synthetic datasets. In the first dataset, Naive CP achieves over-coverage, while in the second, Naive CP undercovers the response. For each dataset, we apply PCP using weights shifted by $\delta$, as in (7). We refer to Appendix D.3 for the full details about this experimental setup. The coverage rates of PCP for different values of $\delta$ are shown in Figure 2. This figure indicates that when Naive CP over-covers, PCP achieves valid coverage for $\delta\geq 0$ or $\delta<-\frac{{W}_{n+1}}{n+1}\approx-1$. However, when Naive CP undercovers, $\delta$ must lie within $\left(-\frac{{W}_{n+1}}{n+1},0\right)$ for PCP to achieve the nominal coverage rate. This result is connected to the MEPS experiment from Figure 1 in which Naive CP and PCP undercover the response, indicating that the weight error does not fall inside the interval. To conclude, the empirical regions of $\delta$ in which PCP is valid that we observed in this experiment align with the theoretical bounds from Theorem 2.
@@
We assume that the errors are bounded, i.e., $\delta_{i}\in[{\delta_{\text{min}}},{\delta_{\text{max}}}]$ for some ${\delta_{\text{min}}},{\delta_{\text{max}}}\in\mathbb{R}$ and $\tilde{\delta}_{i}=(\delta_{i}-{\delta_{\text{min}}})/({\delta_{\text{max}}}-{\delta_{\text{min}}})$ are the normalized errors. We denote the XOR operator by $(a\text{ XOR }b)=((\text{not }a)\text{ and }b)\text{ or }(a\text{ and }(\text{not }b))$, and the NXOR operator by $(a\text{ NXOR }b)=(\text{not }(a\text{ XOR }b))$. We also denote by $k^{\texttt{WCP}}$ the index of the score corresponding to the threshold $Q^{\texttt{WCP}}$ from (5), generated by WCP with true weights: $k^{\texttt{WCP}}:=\min\left\{k:\sum_{i=1}^{k}w_{i}/{W}_{n+1}\geq 1-\alpha\right\}$. With these notations in mind, we now derive the conditions under which PCP applied with weights of a general error achieves valid coverage.
@@
    $\mathbb{P}\left(Q^{\texttt{CP}}>Q^{\texttt{WCP}}\text{, }\left({\bar{\delta}}\leq\frac{\tilde{\Delta}_{n+1}{W}_{k^{\texttt{WCP}}}-\tilde{\Delta}_{k^{\texttt{WCP}}}{W}_{n+1}}{{W}_{n+1}k^{\texttt{WCP}}-(n+1){W}_{k^{\texttt{WCP}}}}\text{ NXOR }{\delta_{\text{min}}}>-\frac{{\delta_{\text{max}}}\tilde{\Delta}_{n+1}+{W}_{n+1}}{n+1-\tilde{\Delta}_{n+1}}\right)\right)\geq 1-\varepsilon$,
@@
    $\mathbb{P}\left(Q^{\texttt{CP}}<Q^{\texttt{WCP}}\text{, }\left({\bar{\delta}}\leq\frac{\tilde{\Delta}_{n+1}{W}_{k^{\texttt{WCP}}}-\tilde{\Delta}_{k^{\texttt{WCP}}}{W}_{n+1}}{{W}_{n+1}k^{\texttt{WCP}}-(n+1){W}_{k^{\texttt{WCP}}}}\text{ XOR }{\delta_{\text{min}}}>-\frac{{\delta_{\text{max}}}\tilde{\Delta}_{n+1}+{W}_{n+1}}{n+1-\tilde{\Delta}_{n+1}}\right)\right)\geq 1-\varepsilon$,
@@
Above, $\tilde{\Delta}_{k}:=\sum_{i=1}^{k}\tilde{\delta}_{i}$, and ${\bar{\delta}}:={\delta_{\text{min}}}/({\delta_{\text{max}}}-{\delta_{\text{min}}})$. Then, the prediction set ${C}^{\texttt{PCP}}(X^{\text{test}})$ constructed by PCP with weights shifted by a general error, as in (9), attains the following coverage rate:
@@ 3.2 Case 2: when the PI explains the label
In this section, we introduce uncertain imputation (UI), a novel and different approach to address corrupted labels, which, in contrast to PCP, does not require access to the conditional corruption probabilities. With UI, however, we use the PI, which is always observed, to impute the corrupted labels with an uncertain version of them, and show that this procedure achieves a valid coverage rate under several assumptions.
@@
We define $\mathcal{E}(z)$ as a reference set of holdout errors conditional on $z$, i.e., $\mathcal{E}(z):=\{E_{i}:i\in\mathcal{I}_{3},Z_{i}=z\}$ and denote by $E(z)$ a random variable drawn uniformly from $\mathcal{E}(z)$. We impute the corrupted labels using this set:
@@
Next, we compute the non-conformity scores of the imputed calibration set: $\bar{S}_{i}=\mathcal{S}({X}_{i},\bar{Y}_{i};\hat{f}),\forall i\in\mathcal{I}_{2}$, and define the threshold using these scores:
@@
$C^{\texttt{UI}}(X^{\text{test}})=\left\{y:\mathcal{S}(X^{\text{test}},y,\hat{f})\leq Q^{\texttt{UI}}\right\}.$ (14)
We summarize this procedure in Algorithm 2 in Appendix C.3. We now show that UI achieves a valid marginal coverage rate if (1) $\hat{g}$ is sufficiently accurate; and (2) the set $C^{\texttt{UI}}$ contains all peaks of the distribution of $Y\mid X,Z$, as formulated next.
@@
Suppose that $\{({X}_{i},{Y}_{i},\tilde{Y}_{i},Z_{i},M_{i})\}_{i=1}^{n+1}$ are i.i.d., and $(X,Y)\perp\!\!\!\perp M\mid Z$. Denote by $C^{\texttt{UI}}(x)=[a(x),b(x)]$ the prediction set constructed by UI that draws errors from the true distribution of $E\mid Z$. Suppose that $Y$ follows the model:
$Y=g^{*}(X,Z)+\varepsilon,$ (15)
where $\varepsilon$ is drawn from a distribution $P_{E^{*}}$ and $\varepsilon\perp\!\!\!\perp X\mid Z$. Further, suppose that
@@
    The model $\hat{g}$ is sufficiently accurate so that there exists a residual error $R^{\text{test}}$ satisfying: (a) $\hat{g}(X^{\text{test}},Z^{\text{test}})=g^{*}(X^{\text{test}},Z^{\text{test}})+R^{\text{test}}$, and (b) $R^{\text{test}}\perp\!\!\!\perp(g^{*}(X^{\text{test}},Z^{\text{test}}),C^{\texttt{UI}}(X^{\text{test}}))\mid Z^{\text{test}}$.
@@
      $\forall v>0:f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(b(x)+v)\geq f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(b(x)-v)$, and
@@
      $\forall v>0:f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(a(x)-v)\geq f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(a(x)+v)$.
Then,
$\mathbb{P}(Y^{\text{test}}\in C^{\texttt{UI}}(X^{\text{test}}))\geq 1-\alpha.$ (16)
In Appendix A.2, we show that under the assumptions of Theorem 4, the imputed labels $\bar{Y}$ have the same distribution as the true labels $Y$ with an added symmetric noise. For example, this noise can be thought of as a normally distributed error with mean zero. The label-noise theory from [13] suggests that using CP with labels affected by symmetric noise yields conservative prediction sets, which guarantees the validity of UI. We remark that in practical applications, if $Z$ is continuous or high-dimensional, the reference set $\mathcal{E}(z)$ might be empty or too small. To alleviate this, we recommend employing the clustering techniques described in Appendix C.3.1.
@@ 3.3 Triply robust conformal prediction with privileged information
The triply robust prediction set is guaranteed to achieve the nominal coverage rate if the requirements of at least one of the methods are satisfied, as stated next.
@@
Suppose that $\{({X}_{i},{Y}_{i},\tilde{Y}_{i},Z_{i},M_{i})\}_{i=1}^{n+1}$ are exchangeable, and $(X,Y)\perp\!\!\!\perp M\mid Z$. Further, suppose that at least one of the following is satisfied: (1) The model $\hat{f}$ is sufficiently accurate so that the scores $\{\mathcal{S}(X_{i},\tilde{Y}_{i};\hat{f}):i=1,...,n,M_{i}=0\}\cup\{\mathcal{S}(X^{\text{test}},Y^{\text{test}};\hat{f})\}$ are exchangeable; (2) The assumptions of Theorem 1 hold; (3) The assumptions of Theorem 4 hold. Then, the prediction set $C^{\texttt{TriplyRobust}}(X^{\textup{test}})$ from (17) achieves the desired coverage rate:
@@ 4 Experiments
In this section, we quantify the effectiveness of our proposed techniques through three experiments. In all experiments, the dataset is randomly split into four distinct subsets: training, validation, calibration, and test sets. For the UI method, the original calibration set is further partitioned into a reference set and a calibration set. The training set is used for fitting a predictive model, and the validation set is used for early stopping. The model is then calibrated using the calibration data, and the performance is evaluated on the test set. Moreover, we use the CQR [20] non-conformity scores, aiming to obtain $1-\alpha=90\%$ coverage rate. The full details regarding the training methodology, datasets, corruption techniques, and the overall experimental protocol are given in Appendix E. We further demonstrate our proposals on a causal inference task using the NSLM dataset [21] in Appendix F.3.
@@ 4.1 Synthetic experiment: the robustness of TriplyRobust
We demonstrate the robustness of TriplyRobust by combining degenerate and oracle variants of each underlying component: quantile regression (QR), PCP, and UI. We generate a synthetic dataset with 10-dimensional inputs $X$, 3-dimensional privileged information $Z$, and a continuous response $Y$ that is a function of $X$ and $Z$. The dataset is artificially corrupted by removing labels, such that labels with greater uncertainty have a higher probability of being removed.. The degenerate variant of QR outputs the trivial prediction set $\{0\}$, while the oracle version uses the true conditional quantiles of $Y\mid X$ to achieve the desired conditional coverage. Similarly, the degenerate version of PCP computes weights using half the true corruption probabilities. In contrast, its oracle counterpart employs the correct probabilities. The trivial imputation method assigns missing labels the value 00, while the oracle method draws labels from the true conditional distribution of $Y\mid Z$. Figure 3 shows that when all three components are degenerate, TriplyRobust achieves a coverage rate lower than the nominal level. However, when at least one technique is oracle-based, TriplyRobust produces valid uncertainty estimates. This experiment reveals that the coverage rate achieved by TriplyRobust is not overly conservative, despite being a union of three intervals.
Figure 3: The coverage rate achieved by TriplyRobust with “degenerate” or “oracle” underlying models; target coverage $1-\alpha=90\%$. The coverage is evaluated over 30 independent trials.

### v3

@@ 3.1.2 Is PCP robust to inaccurate weights?
The following robustness analysis of PCP is divided into two parts. First, we consider a case where the inaccurate weights $\{\tilde{w}_{i}\}_{i=1}^{n}$ are shifted by a constant error $\delta\in\mathbb{R}$ from the true weights for all samples. In the second, we extend the analysis to a more general setting where the error varies across samples. We remark that the theory developed in this section also applies to WCP, as detailed in Appendix A.1. We begin by examining the case of constant error, formulated as:
@@
We remark that, in this analysis, we do not consider the sign of the weights and allow them to become negative. We denote the sum of the true weights by ${W}_{k}:=\sum_{j=1}^{k}w_{j}$ and recall that $Q^{\texttt{WCP}}$ from equation 4 is the threshold constructed by WCP using the true weights. We also denote by $Q^{\texttt{CP}}$ the threshold generated by Naive CP, which is CP from Section 2.1 applied using only the uncorrupted data; see Appendix A.1.1 for more details. These notations set the ground for the conditions required for PCP to achieve the desired coverage rate.
@@
Suppose that the assumptions of Theorem 1 hold. Further, suppose that at least one of the following holds: (1) $\mathbb{P}\left(Q^{\texttt{CP}}>Q^{\texttt{WCP}}\right)\geq 1-\varepsilon$ and $\delta\geq 0$; (2) $\mathbb{P}\left(Q^{\texttt{CP}}>Q^{\texttt{WCP}},\delta<-\frac{{W}_{n+1}}{n+1}\right)\geq 1-\varepsilon$; (3) $\mathbb{P}\left(Q^{\texttt{CP}}<Q^{\texttt{WCP}},\delta>-\frac{{W}_{n+1}}{n+1}\right)\geq 1-\varepsilon$ and $\delta\leq 0$; (4) $\mathbb{P}\left(Q^{\texttt{CP}}=Q^{\texttt{WCP}}\right)\geq 1-\varepsilon$. Then, the prediction set ${C}^{\texttt{PCP}}(X^{\text{test}})$ constructed by PCP with weights shifted by $\delta$, as in equation 6, satisfies:
@@
We demonstrate Theorem 2 on two synthetic datasets. In the first dataset, Naive CP achieves over-coverage, while in the second, Naive CP undercovers the response. For each dataset, we apply PCP using weights shifted by $\delta$, as in equation 6. We refer to Appendix D.3 for the full details about this experimental setup. The coverage rates of PCP for different values of $\delta$ are shown in Figure 2. This figure indicates that when Naive CP over-covers, PCP achieves valid coverage for $\delta\geq 0$ or $\delta<-\frac{{W}_{n+1}}{n+1}\approx-1$. However, when Naive CP undercovers, $\delta$ must lie within $\left(-\frac{{W}_{n+1}}{n+1},0\right)$ for PCP to achieve the nominal coverage rate. This result is connected to the MEPS experiment from Figure 1 in which Naive CP and PCP undercover the response, indicating that the weight error does not fall inside the interval. In conclusion, the empirical regions of $\delta$ in which PCP is valid that we observed in this experiment align with the theoretical bounds from Theorem 2.
@@
[…]
@@
    $\mathbb{P}\left(Q^{\texttt{CP}}<Q^{\texttt{WCP}}\text{, }\left({\bar{\delta}}\leq\frac{\tilde{\Delta}_{n+1}{W}_{k^{\texttt{WCP}}}-\tilde{\Delta}_{k^{\texttt{WCP}}}{W}_{n+1}}{{W}_{n+1}k^{\texttt{WCP}}-(n+1){W}_{k^{\texttt{WCP}}}}\text{ NXOR }{\delta_{\text{min}}}>-\frac{{\delta_{\text{max}}}\tilde{\Delta}_{n+1}+{W}_{n+1}}{n+1-\tilde{\Delta}_{n+1}}\right)\right)\geq 1-\varepsilon$,
@@
    $\mathbb{P}\left(Q^{\texttt{CP}}>Q^{\texttt{WCP}}\text{, }\left({\bar{\delta}}\leq\frac{\tilde{\Delta}_{n+1}{W}_{k^{\texttt{WCP}}}-\tilde{\Delta}_{k^{\texttt{WCP}}}{W}_{n+1}}{{W}_{n+1}k^{\texttt{WCP}}-(n+1){W}_{k^{\texttt{WCP}}}}\text{ XOR }{\delta_{\text{min}}}>-\frac{{\delta_{\text{max}}}\tilde{\Delta}_{n+1}+{W}_{n+1}}{n+1-\tilde{\Delta}_{n+1}}\right)\right)\geq 1-\varepsilon$,
@@
Above, $\tilde{\Delta}_{k}:=\sum_{i=1}^{k}\tilde{\delta}_{i}$, and ${\bar{\delta}}:={\delta_{\text{min}}}/({\delta_{\text{max}}}-{\delta_{\text{min}}})$. Then, the prediction set ${C}^{\texttt{PCP}}(X^{\text{test}})$ constructed by PCP with weights shifted by a general error, as in equation 8, attains the following coverage rate:
@@ 3.2 Case 2: when the PI explains the label
In this section, we introduce uncertain imputation (UI), a novel and different approach to address corrupted labels, which, in contrast to PCP, does not require access to the conditional corruption probabilities. With UI, however, we use the PI, which is always observed, to impute the corrupted labels with an uncertain version of them, and show that this procedure achieves a valid coverage rate under several assumptions. This way, UI can obtain valid coverage even when PCP does not. Intuitively, UI is more useful than PCP when $Y$ is relatively easy to predict given $X,Z$.
@@
We define $\mathcal{E}(z)$ as a reference set of holdout errors conditional on $z$, i.e., $\mathcal{E}(z):=\{E_{i}:i\in\mathcal{I}_{3},Z_{i}=z,M_{i}=0\}$ and denote by $E(z)$ a random variable drawn uniformly from $\mathcal{E}(z)$. We impute the corrupted labels using this set, and denote the imputed label by $\bar{Y}_{i}$:
@@
Next, we compute the non-conformity scores of the imputed calibration set, denoted by $\bar{S}_{i}$: $\bar{S}_{i}=\mathcal{S}({X}_{i},\bar{Y}_{i};\hat{f}),\forall i\in\mathcal{I}_{2}$, and define the threshold using these scores:
@@
Finally, for a new input data $X^{\text{test}}$, we construct the prediction set for $Y^{\text{test}}$ as: $C^{\texttt{UI}}(X^{\text{test}})=\left\{y:\mathcal{S}(X^{\text{test}},y,\hat{f})\leq Q^{\texttt{UI}}\right\}$. We summarize this procedure in Algorithm 2 in Appendix C.3. We now show that UI achieves a valid marginal coverage rate if (1) $\hat{g}$ is sufficiently accurate; and (2) the set $C^{\texttt{UI}}$ contains all peaks of the distribution of $Y\mid X,Z$, as formulated next.
@@
Suppose that $\{({X}_{i},{Y}_{i},\tilde{Y}_{i},Z_{i},M_{i})\}_{i=1}^{n+1}$ are i.i.d., and $(X,Y)\perp\!\!\!\perp M\mid Z$. Denote by $C^{\texttt{UI}}(x)=[a(x),b(x)]$ the prediction set constructed by UI that draws errors from the true distribution of $E\mid Z$. Suppose that $Y$ follows the model: $Y=g^{*}(X,Z)+\varepsilon$, where $\varepsilon$ is drawn from a distribution $P_{E^{*}}$ and $\varepsilon\perp\!\!\!\perp X\mid Z$. Further, suppose that
@@
    $\hat{g}$ is sufficiently accurate so that there exists a residual $R^{\text{test}}$ satisfying: (a) $\hat{g}(X^{\text{test}},Z^{\text{test}})=g^{*}(X^{\text{test}},Z^{\text{test}})+R^{\text{test}}$, and (b) $R^{\text{test}}\perp\!\!\!\perp(g^{*}(X^{\text{test}},Z^{\text{test}}),C^{\texttt{UI}}(X^{\text{test}}))\mid Z^{\text{test}}$.
@@
      $\forall v>0:f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(b(x)+v)\leq f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(b(x)-v)$, and
@@
      $\forall v>0:f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(a(x)-v)\leq f_{Y^{\text{test}}\mid X^{\text{test}}=x,Z^{\text{test}}=z}(a(x)+v)$.
Then, $\mathbb{P}(Y^{\text{test}}\in C^{\texttt{UI}}(X^{\text{test}}))\geq 1-\alpha$.
The proof is given in Appendix A.2. We pause here to clarify the assumptions of Theorem 4. The first assumption requires that the residual errors are independent of the prediction of $\hat{g}$ and of $C^{\texttt{UI}}$ given the PI $Z$. Notably, this assumption does not restrict the distribution of the errors or their magnitude and holds even under substantial errors. Intuitively, this means that the PI serves as a good proxy for $Y$. For example, in medical imaging, a pathologist’s diagnostic report can act as privileged information that predicts tissue diagnosis. In this case, we require that the predictor that estimates $Y$ from the PI is accurate up to an error that is independent of $g^{*},C$ conditional on $Z$. The second assumption of Theorem 4 states that the distribution of $Y\mid X,Z$ is concentrated in the predicted interval, and that the density at $Y=y$ decreases as we move away from this interval. Importantly, this assumption does not limit the number of peaks in the distribution of $Y\mid X,Z$, as long as all such peaks are inside the interval. Moreover, since the PI is strongly indicative of the label in our framework, the distribution of $Y\mid X,Z$ is expected to be relatively simple in practice. Overall, this assumption is relatively mild, particularly for prediction intervals that aim to achieve a high coverage rate. In Appendix F.5 demonstrate in experiments that UI can obtain valid coverage even when the independence requirements of this theorem are not exactly satisfied. We remark that in practical applications, if $Z$ is continuous or high-dimensional, the reference set $\mathcal{E}(z)$ might be empty or too small. To alleviate this, we recommend employing the clustering techniques described in Appendix C.3.1.
@@ 3.3 Triply robust conformal prediction with privileged information
This approach achieves the nominal coverage rate if the assumptions of one of the methods hold:
@@
[…]
@@ 4 Experiments
In this section, we quantify the effectiveness of our proposed techniques through three experiments. In all experiments, the dataset is randomly split into four distinct subsets: training, validation, calibration, and test sets. For the UI method, the original calibration set is further partitioned into a reference set and a calibration set. The training set is used for fitting a predictive model, and the validation set is used for early stopping. The model is then calibrated using the calibration data, and the performance is evaluated on the test set over 30 random data splits. Moreover, we use the CQR (Romano et al., 2019) non-conformity scores, aiming to obtain $1-\alpha=90\%$ coverage rate. The full details regarding the training methodology, datasets, corruption techniques, and the overall experimental protocol are given in Appendix E. We further demonstrate our proposals on a causal inference task using the NSLM dataset (Yeager et al., 2019) in Appendix F.3.
@@ 4.1 Synthetic experiment: the robustness of TriplyRobust
We demonstrate the robustness of TriplyRobust by combining degenerate and oracle variants of each underlying component: quantile regression (QR), PCP, and UI. We generate a synthetic dataset with 10-dimensional inputs $X$, 3-dimensional privileged information $Z$, and a continuous response $Y$ that is a function of $X$ and $Z$. The dataset is artificially corrupted by removing labels, such that labels with greater uncertainty have a higher probability of being removed. The degenerate variant of QR outputs the trivial prediction set $\{0\}$, while the oracle version uses the true conditional quantiles of $Y\mid X$ to achieve the desired conditional coverage. Similarly, the degenerate version of PCP computes weights using half the true corruption probabilities. In contrast, its oracle counterpart employs the correct probabilities. The trivial imputation method assigns missing labels the value 0, while the oracle method draws labels from the true conditional distribution of $Y\mid Z$. Figure 3 shows that when all three components are degenerate, TriplyRobust achieves a coverage rate lower than the nominal level. However, when at least one technique is oracle-based, TriplyRobust produces valid uncertainty estimates. This experiment reveals that the coverage rate achieved by TriplyRobust is not overly conservative, despite being a union of three intervals.
Figure 3: The coverage achieved by TriplyRobust with “degenerate” or “oracle” models.